OpenStudy (anonymous):

Solve for y in the following system of equations. -8x - 8y = -64 8x - 8y = -96 anyone PLEASE im in a class and i really need help and if i dont get threw this than i dont graduate. its not that i didnt try in the class the teacher was a ss and i cant leran with that.

OpenStudy (anonymous):

Treat this like a simple addition/subtraction problem -8x+(-8y)=-64 + 8x+(-8y)=-96 -------------- 0x+(-16y)=-160 -16y=-160 y= 10